Find the equation of the solution to dy/dx = x^(5) * y through the point (x;y)=(1;2)



1
Expert's answer
2022-03-14T19:57:26-0400

"\frac{dy}{dx}=x⁵y," "(x,y)=(1,2)"


"\frac{dy}{y}=x⁵dx"


Integrating both sides, we have;


"lny=\frac{x⁶}{6}+c"


Recall that y(1)=2


"ln2=\frac{1}{6}+c"


"c=ln2-\frac{1}{6}"


"c=0.53"


"lny=\frac{x⁶}{6}+0.53"


"=>y=e^{\frac{x⁶}{6}+0.53}"


"=>y=e^{\frac{x⁶}{6}}e^{0.53}"


"=>y=1.70e^{\frac{x⁶}{6}}"